OpenStudy (anonymous):

find solution not factors: x^2+30x+216=0

sam (.sam.):

x=-18,-12

sam (.sam.):

(x+18)(x+12)=0 x=-18,-12

OpenStudy (anonymous):

how did u do that

sam (.sam.):

Factor

sam (.sam.):